# Noé Arnold, Juan Beccuti, Thunj Chantramonklasri, Matthias Hafner, Nicolas Oderbolz Center for Cryptoeconomics

### v1

@@ IV-B1 Baseline Model
We first compare the equilibrium staking outcomes of the calibrated baseline model with the equilibrium outcomes of the calibrated model under the alternative issuance schedule $y^{\prime}_{i}(D)$. Table I presents the total amount of ETH staked and the corresponding profits for each type of ETH holder under various issuance schedules, as well as the relative changes in these values resulting from the change in the issuance schedule.
TABLE I: Nash Equilibria of Baseline Model
Type Num $y_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}}$ $y^{\prime}_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}\cdot(1+k\cdot D)}$ $\Delta$
Deps Ratio Profita Deps Ratio Profita Deps Profit
e 25K 0.9M 2.7% 0.8% 0.74M 2.97% 0.2% -17.8% -75%
t 200K 18M 54.1% 1.8% 14.3M 57.5% 1.6% -20.6% -11.1%
r 925K 14.4M 43.2% 0.95% 9.8M 39.53% 0.78% -31.5% -17.9%
Total - 33.3M - - 24.84M - - -25.4% -
a Profit per ETH staked.
@@
The first observation to make is that a change in the issuance schedule results in an approximate 25% reduction in the total amount of staked ETH. The relative decrease is most pronounced for the ETH holder type Retailer, followed by Techie and Expert. Consequently, although the overall staking supply declines across all types, the market share of solo staking and staking via dSSPs increases. This result may be explained by Experts facing high marginal costs when solo staking, making them less sensitive to changes in the issuance schedule, as well as Techies having access to additional external DeFi yields, again making them less sensitive to changes in the issuance schedule. Despite a substantial reduction in their equilibrium staking supply, all ETH holder types experience lower staking profits per ETH staked under the adjusted issuance schedule. This decline is particularly pronounced for the Expert type, who uniquely must cover their fixed costs and higher average costs with a reduced revenue base.
IV-B2 Extended Model - Expected MEV Revenues
@@
Next, we repeat the above analysis for the extended version of the model, where the type Expert is assumed to be risk-averse and as such discounts future MEV revenues. Table II again reports the resulting equilibrium outcomes under the two issuance schedules.
@@
TABLE II: Nash Equilibria of Extended Model with Discounted MEV Revenues
Type Num $y_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}}$ $y^{\prime}_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}\cdot(1+k\cdot D)}$ $\Delta$
Deps Ratio Profita Deps Ratio Profita Deps Profit
e 25K 0.77M 2.3% 0.6% 0.58M 2.3% -0.01% -24.6% -100%
t 200K 18M 54.3% 1.8% 14.3M 57.8% 1.6% -20.6% -11.1%
r 925K 14.4M 43.4% 0.95% 9.9M 39.85% 0.78% -31.25% -17.9%
Total - 33.2M - - 24.76M - - -25.4% -
a Profit per ETH staked.
@@
When incorporating the variance in MEV rewards for a risk-averse and small solo staker of type Expert, their relative reduction in staking supply becomes more pronounced compared to the baseline model. As outlined in Section IV-A, this outcome is driven by the following dynamic: if agents expect to have access to lower expected MEV revenues, they become more sensitive to changes in the issuance schedule. As a result, a change in the issuance schedule no longer leads to an increase in the market share of solo stakers. At the same time the staking outcomes for the remain staking solutions and ETH holder types remain virtually unchanged relative to the baseline model. Since Expert solo stakers represent only a small fraction of the overall population, their heightened sensitivity to issuance changes has little impact on Techies and Retailers. Additionally, the negative effect of issuance reductions on the expected profits per staked ETH for solo stakers is further exacerbated, with expected profits even turning negative.
IV-B3 Extended Model - Inattentive Agents
@@
Now, we incorporate to the model with the higher variability in MEV rewards for solo stakers, the differentiation of the ETH holder type Retailer into (1) institutions and (2) an inattentive retail investor, as described in Section III-B2. Table III reports the corresponding results.
@@
TABLE III: Nash Equilibria of Extended Model with Inattentive Agents
Type Num $y_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}}$ $y^{\prime}_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}\cdot(1+k\cdot D)}$ $\Delta$
Deps Ratio Profita Deps Ratio Profita Deps Profit
e 25K 0.78M 2.4% 0.6% 0.57M 2.2% -0.0% -26.9% -100%
t 200K 18.1M 54.7% 1.8% 13.7M 53.3% 1.6% -24.3% -11.1%
i 300K 7.7M 23.3% 0.95% 4.9M 19.1% 0.76% -36.4% -20.0%
r NA 6.5M 19.6% NA 6.5M 25.3% NA 0% NA
Total - 33.1M - - 25.7M - - -22.4% -
a Profit per ETH staked.
@@
In the presence of the inattentive ETH holder type—whose staking supply remains constant across both issuance schedules—the relative reduction in the overall staking supply is somewhat lower than in the baseline model. Meanwhile, the ETH holder types that do adjust their staking supply in response to the issuance schedule experience a decline in their market shares under the adjusted schedule. Again this is consistent with the intuitions developed in Section IV-A. When one ETH holder type does not modify its staking supply in response to an issuance reduction, the adjustments made by the other types tend to be more pronounced. The solo staking type Expert still experiences a larger reduction in their staking supply than the Techie type, who continues to benefit from external DeFi yields, making their equilibrium staking supply less sensitive to changes in protocol issuance. Consequently, the decline in staking profits per ETH staked is even greater for the Expert type.
IV-B4 Extended Model - Intermediary dSSP with Market Power
@@
Finally, we conduct a comparative analysis of equilibrium outcomes under the assumption of an intermediary dSSP, similar to Lido, that acts as a platform between ETH holders and SSPs. Importantly, the fee set by the dSSP is endogenous to the model and therefore subject to the strategic response of the dSSP to a change in the issuance schedule. In solving the model, we make some additional assumptions and simplifications relative to the model presented in Section III-B3:
@@
Table IV reports the results of the comparative analysis under these assumptions, in particular the fees set by the dSSP and the distribution of stakes across different ETH holder types and staking methods.
TABLE IV: Nash Equilibria of Extended Model with Intermediary dSSP
Type Num $y_{i}(D);f_{d}=10\%$ $y^{\prime}_{i}(D);f_{d}=13\%$ $\Delta$
Deps Ratio Deps Ratio Deps
ss 25K 0.78M 2.4% 0.57M 2.2% -26.9%
t 200K 18M 54.5% 13.4M 52.5% -25.5%
i 300K 7.7M 23.3% 5M 19.6% -35%
r NA 6.5M 19.7% 6.5M 25.5% 0%
Total 33M 25.5M -22.7%
a Dep = Deposits, Ratio = Staking Ratio, $\Delta$ = Change in Deposits
@@
V Discussion

### v2

@@ 4.7.3 Comparative Analysis
Table 3 presents the total amount of ETH staked and the corresponding profits for each type of ETH holder under various issuance schedules, as well as the relative changes in these values resulting from the change in the issuance schedule.
Table 3: Nash equilibria of baseline model
$y_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}}$ $y^{\prime}_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}\cdot(1+k\cdot D)}$ $\Delta$
Types Num Deps Ratio Profita Deps Ratio Profita Deps Profit
e $25K$ $0.9M$ $2.7\%$ $0.8\%$ $0.74M$ $2.97\%$ $0.2\%$ $-17.8\%$ $-75\%$
t $200K$ $18M$ $54.1\%$ $1.8\%$ $14.3M$ $57.5\%$ $1.6\%$ $-20.6\%$ $-11.1\%$
r $925K$ $14.4M$ $43.2\%$ $0.95\%$ $9.8M$ $39.53\%$ $0.78\%$ $-31.5\%$ $-17.9\%$
Total $33.3M$ $24.84M$ $-25.4\%$
a Profit per ETH staked
@@
The first observation is that a change in the issuance schedule results in an approximate 25% reduction in the total amount of staked ETH. The relative decrease is most pronounced for the ETH holder type Retailer, followed by Techie and Expert. Consequently, although the overall staking supply declines across all types, the market share of solo staking and staking via dSSPs increases. Based on the intuitions developed in Section 4.7.1, this result can be explained by Experts facing high marginal costs when solo staking, making them less sensitive to changes in the issuance schedule, as well as Techies having access to additional external DeFi yields, again making them less sensitive to changes in the issuance schedule. Despite a substantial reduction in their equilibrium staking supply, all ETH holder types experience lower staking profits per ETH staked under the adjusted issuance schedule. This decline is particularly pronounced for the Expert type, who uniquely must cover their fixed costs and higher average costs with a reduced revenue base.
Expected MEV Revenues
@@
Next, we repeat the above analysis for the extended version of the model, where the type Expert is assumed to be risk-averse and as such discounts future MEV revenues. Table 4 again reports the resulting equilibrium outcomes under the two issuance schedules.
@@
Table 4: Nash equilibria of extended model with discounted MEV revenues
$y_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}}$ $y^{\prime}_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}\cdot(1+k\cdot D)}$ $\Delta$
Types Num Deps Ratio Profita Deps Ratio Profita Deps Profit
e $25K$ $0.77M$ $2.3\%$ $0.6\%$ $0.58M$ $2.3\%$ $0.0\%$ $-24.6\%$ $-100\%$
t $200K$ $18M$ $54.3\%$ $1.8\%$ $14.3M$ $57.8\%$ $1.6\%$ $-20.6\%$ $-11.1\%$
r $925K$ $14.4M$ $43.4\%$ $0.95\%$ $9.9M$ $39.85\%$ $0.78\%$ $-31.25\%$ $-17.9\%$
Total $33.2M$ $24.76M$ $-25.4\%$
a Profit per ETH staked
@@
When incorporating the variance in MEV rewards for a risk-averse and small solo staker of type Expert, their relative reduction in staking supply becomes more pronounced compared to the baseline model. As outlined in Section 4.7.1, this outcome is driven by the following dynamic: if agents expect to have access to lower expected MEV revenues, they become more sensitive to changes in the issuance schedule. As a result, a change in the issuance schedule no longer leads to an increase in the market share of solo stakers. At the same time the staking outcomes for the remain staking solutions and ETH holder types remain virtually unchanged relative to the baseline model. Since Expert solo stakers represent only a small fraction of the overall population, their heightened sensitivity to issuance changes has little impact on Techies and Retailers. Additionally, the negative effect of issuance reductions on the expected profits per staked ETH for solo stakers is further exacerbated, with expected profits even turning negative.
Inattentive Agents
@@
Now, we incorporate to the model with the higher variability in MEV rewards for solo stakers, the differentiation of the ETH holder type Retailer into (1) institutions and (2) an inattentive retail investor, as described in Section 4.6.2. Table 5 reports the corresponding results.
@@
Table 5: Nash equilibria of extended model with inattentive agents
$y_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}}$ $y^{\prime}_{i}(D)=\frac{2.6\cdot 64}{\sqrt{D}\cdot(1+k\cdot D)}$ $\Delta$
Types Num Deps Ratio Profita Deps Ratio Profita Deps Profit
e $25K$ $0.78M$ $2.4\%$ $0.6\%$ $0.57M$ $2.2\%$ $0.0\%$ $-26.9\%$ $-100\%$
t $200K$ $18.1M$ $54.7\%$ $1.8\%$ $13.7M$ $53.3\%$ $1.6\%$ $-24.3\%$ $-11.1\%$
i $300K$ $7.7M$ $23.3\%$ $0.95\%$ $4.9M$ $19.1\%$ $0.76\%$ $-36.4\%$ $-20.0\%$
r NA $6.5M$ $19.6\%$ NA $6.5M$ $25.3\%$ NA $0\%$ NA
Total $33.1M$ $25.7M$ $-22.4\%$
a Profit per ETH staked
@@
In the presence of the inattentive ETH holder type—whose staking supply remains constant across both issuance schedules—the relative reduction in the overall staking supply is somewhat lower than in the baseline model. Meanwhile, the ETH holder types that do adjust their staking supply in response to the issuance schedule experience a decline in their market shares under the adjusted schedule. Again this is consistent with the intuitions developed in Section 4.7.1. When one ETH holder type does not modify its staking supply in response to an issuance reduction, the adjustments made by the other types tend to be more pronounced. The solo staking type Expert still experiences a larger reduction in their staking supply than the Techie type, who continues to benefit from external DeFi yields, making their equilibrium staking supply less sensitive to changes in protocol issuance. Consequently, the decline in staking profits per ETH staked is even greater for the Expert type.
Intermediary dSSP with Market Power
@@
Finally, we conduct a comparative analysis of equilibrium outcomes under the assumption of an intermediary dSSP, similar to Lido, that acts as a platform between ETH holders and SSPs. Importantly, the fee set by the dSSP is endogenous to the model and therefore subject to the strategic response of the dSSP to a change in the issuance schedule. In solving the model, we make some additional assumptions and simplifications relative to the model presented in Section 4.6.3:
@@
Table 6 reports the results of the comparative analysis under these assumptions, in particular the fees set by the dSSP and the distribution of stakes across different ETH holder types and staking methods.
Table 6: Nash equilibria of extended model with intermediary dSSP
$y_{i}(D);f_{d}=10\%$ $y^{\prime}_{i}(D);f_{d}=13\%$ $\Delta$
Types Num Deps Ratio Deps Ratio Deps
ss $25K$ $0.78M$ $2.4\%$ $0.57M$ $2.2\%$ $-26.9\%$
t $200K$ $18M$ $54.5\%$ $13.4M$ $52.5\%$ $-25.5\%$
i $300K$ $7.7M$ $23.3\%$ $5M$ $19.6\%$ $-35\%$
r NA $6.5M$ $19.7\%$ $6.5M$ $25.5\%$ $0\%$
Total $33M$ $25.5M$ $-22.7\%$
@@
5 Empirical Analysis
In the following, we employ an instrumental variables (IV) approach to obtain unbiased estimates of the slope of the supply curves across different staking methods.
5.1 Estimation Approach
To address endogeneity concerns in the estimates of staking supply elasticity, we employ an instrumental variable (IV) approach. Endogeneity arises in this context because any observed combination of staking yields and staking amounts reflects a market equilibrium where supply and demand for staking meet. Changes in this equilibrium can result from shifts in either the supply or demand curve. As depicted in Figure 7, the IV approach addresses this issue by identifying instruments that shift the demand for staked ETH without simultaneously affecting the factors that determine the supply curve. This allows us to isolate the slope of the supply curve at the equilibrium point, i.e. the yield elasticity of supply.
Figure 7: Illustration of Instrumental Variable Approach
Consider the following linear model of the staking supply curve:
$D_{jt}=\beta_{0}+\beta_{1}y_{jt}+\gamma^{\prime}W_{t}+\epsilon_{jt},$ (8)
where $D_{jt}$ is the the staking supply at time $t$ for staking method $j$, $y_{jt}$ is the real staking yield at at time $t$ for staking method $j$, and $W_{t}$ a vector of control variables. To obtain a consistent estimate of $\beta_{1}$, we introduce an instrumental variable $Z_{jt}$ that shifts $y_{jt}$ independently of $\epsilon_{jt}$. A valid instrument $Z_{jt}$ must satisfy two key conditions:
  1. 1.
    Instrument relevance ($\mathrm{Cov}(Z_{jt},y_{jt})\neq 0$): The instrument must identify meaningful variation in the independent variable of interest $y_{jt}$. In the literature, this condition is typically argued to be fulfilled if the F-Statistic of the first stage regression, i.e. the regression of $y_{jt}$ on $Z_{jt}$ is greater than 10.
  2. 2.
    Instrument exogeneity ($\mathrm{Cov}(Z_{jt},\epsilon_{jt})=0$): As outlined above, the instrument should be exogenous and the effect of the instrument $Z_{jt}$ on $D_{jt}$ should run solely through its effect on $y_{jt}$. In general, this condition cannot be tested empirically.
Under these conditions, variations in $Z_{i}$ induce exogenous variation in $y_{t}$, which can be used to identify $\beta_{1}$. Formally, we implement a two-stage least squares (2SLS). In the first stage, we regress the endogenous regressor on the instrument:
[…]
obtaining the fitted values $\hat{X}_{i}$. In the second stage, we regress $Y_{i}$ on the fitted values from the first stage:
[…]
yielding the 2SLS estimate $\hat{\beta}_{1}^{IV}$. In the following, we discuss EIP updates and gas fees as potential valid instrumental variables.
5.1.1 EIPs as Instrument
Staked ETH is the way through which Ethereum secures its Proof-of-Stake (PoS) consensus, for which the Ethereum blockchain rewards stakers with ETH. Put differently, demand for staked ETH is the protocol’s demand for security. With this in mind, Ethereum Improvement Proposals (EIPs) involving upgrades that affect how much stakers earn for providing PoS security arguably represent shifts in the demand curve from which one could trace out the staking supply curve.
Regarding the condition of instrument exogeneity, it is not unambiguous that EIP changes do not at all affect the supply curve. For example, part of the Shapella upgrade on April 12th 2023 was introducing the ability for stakers to withdraw their rewards and original stakes. Until then, all rewards were locked on the consensus layer (Beacon Chain). To the extent that future rewards are discounted (due to the time value of money, etc.) the ability to withdraw contemporaneously increases ETH rewards from the protocol, shifting the demand curve to the right. However, it could be argued that would-be stakers anticipating this could be more willing to supply a certain amount of ETH for every yield level, that changes in their expectations would shift the supply curve down/to the right.
Further, The EIP upgrades were all announced long before they occurred, and being mainnet forks, necessarily permanent. Thus, it could be argued that shifts in the supply curve would have been spread out over time, such that the effects captured by a dummy variable for the day of the fork would have captured much more of the direct effect on demand.
A dummy variable was created for each of six relevant EIP forks, starting with the London upgrade on August 5th 2021 through to March 13th 2024’s Dencun upgrade. Changes in PoS rewards captured may be in real or nominal ETH, e.g., EIP-1559 introduced through the London fork changed the fee market structure, separating transaction fees into base and priority fees, the former of which was burned. This effectively increased real ETH rewards for staking and represents the demand curve for staked ETH shifting right. Formally, we estimate the following first stage regression:
$y_{t}=\pi_{0}+\pi_{1}Bellatrix_{t}+\pi_{2}Paris_{t}+\pi_{3}Shapella_{t}+\pi_{4}Dencun_{t}+\pi^{\prime}W_{t}+\nu_{i},$ (11)
where $y_{t}$ represents the log of annualized staking rewards in USD. In the second-stage regression, the predicted value of log USD staking rewards $\hat{y}$ is then used as an explanatory variable:
$D_{t}=\gamma_{0}+\gamma_{1}\hat{y}_{t}+\gamma^{\prime}W_{t}+u_{t}.$ (12)
where $D_{t}$ is the natural log of total quantity staked at time $t$. To control for other factors influencing changes in staking supply decisions, we include additional control variables, e.g., a proxy for the risk-free rate, as well as a dummy for ETH price crashes, to control for the opportunity cost and psychological cost of staking, respectively.
5.1.2 Gas Fees as Instrument
In addition, we explore gas fees as a potential instrument. The causal pathway from gas fees to staking rewards – and subsequently to staking supply – is well-defined and straightforward. Gas fees, on one hand, directly influence the rewards that validators earn from staking. As gas fees increase due to, for example, higher network activity, the effective yield from staking rises. This makes gas fees a relevant instrument, as these fees clearly impact the variable of interest – the staking rewards. On the other hand, gas fees are primarily driven by the demand for and value of transactions on the Ethereum network, which is influenced by factors like network usage, DeFi activity, NFT trading, or other dApp interactions. These factors are largely independent of ETH holders’ staking supply decision, making gas fees an exogenous variable in relation to the staking supply. Thus, using gas fees as an instrument leverages a key component of the Ethereum ecosystem that naturally fluctuates due to external factors (such as user demand) while directly influencing staking rewards but not directly impacting staking supply, making it a robust choice for this analysis. Additionally, the regular fluctuations and public availability of gas fees data enable high-frequency and reliable analysis.
We thus employ the 2SLS approach outline above using gas fees as the instrument. By log-transforming the variables, we can interpret the estimated coefficient $\hat{\beta}_{1}^{IV}$ as the supply elasticity of staking with respect to staking rewards, denominated in USD. Formally, we again estimate the following first stage regression
$log(y_{t})=\pi_{0}+\pi_{1}+log(r_{t})+\nu_{i},$ (13)
where $r_{t}$ represent total gas fees paid in the protocol at time $t-1$. We then specify the second-stage regression as
$log(D_{t})=\gamma_{0}+\gamma_{1}log(\hat{y}_{t})+\gamma^{\prime}W_{t}+u_{t}.$ (14)
5.2 Data
Two main data sets are employed for this analysis. The first is from Rated, an independent provider of node and node operator data, with supplementary data from Dune and Renzo analytics which were kindly shared by fellow grantees. This was the longest set of validator data the study had access to, stemming from the launch of Beacon Chain (i.e., Day 0 or December 1st 2020) through to June 2024. Additional data, e.g., for different PoS L1 real rewards, as well as annualized USD returns from ETH staking, were extracted from Staking Rewards’s API. For ETH, the earliest rewards data was naturally from 1st December 2020. Daily-level data on gas fees were obtained via Dune Analytics. Natural logs of the variables were used, per the derivations above.
5.3 Results
5.3.1 EIPs as an IV
As previously outlined, for an instrument to be valid, it needs to be both relevant to the endogenous variable and uncorrelated with the unobserved error term. While verifying the latter condition is not straightforward, the former is easily tested with an OLS regression of the endogenous variable - the natural log of dollar rewards to staking. With an F-statistic of 428.2, the instruments prove to be collectively relevant to variations in dollar rewards. To the extent the above argument is a compelling one, combined with the encouraging ’pseudo’ first-stage regression, one could implement the EIPs-2SLS method.
Table 7: Testing the relevance of instruments
OLS
Bellatrix $0.0136$
$(0.101)$
Paris $0.1339$
$(0.095)$
Shapella $0.7645^{***}$
$(0.025)$
Dencun $0.5601^{***}$
$(0.041)$
Constant $20.6109^{***}$
$(0.040)$
Observations 423
F-statistic 428.2
R-squared $0.804$
Standard errors in parentheses
*** p$<$0.01, ** p$<$0.05, * p$<$0.1
From Table 8, it can be seen that compared with the OLS case, the coefficient on natural log of dollar rewards from 2SLS is higher, respectively at 0.2086 and 0.4640. Both estimates are positive and significant. This result is non-trivial, as without controlling for the effect of supply shifts, to the extent observed staked amount and yield combinations reflect supply changes more than demand changes (i.e., trace out the demand curve via supply shifts), the coefficient would be negative. Furthermore, the F-statistic of the first-stage regression (i.e., natural log of dollar rewards on the EIP dummies) is 428.2, indicating strong relevance.
The 2SLS estimate above suggests that for every percentage increase in dollar staking reward, staked amount increases by 0.46 percent. The higher value from 2SLS suggests that not accounting for supply-side shocks, the OLS estimator underestimates the dollar reward-elasticity of staking supply. One possible reason for this is that the supply curve shifted to the left during the period of our study. There are numerous possible reasons for this, such as the advent of higher-yielding PoS alternatives motivating substitution away from Ethereum.
Focusing on the remaining 2SLS estimates, the coefficient on the risk-free rate (approximated by the three-month dollar treasury yield) was also positive and significant at the 5% level, at 0.0714. This is in line with the Capital Asset Pricing Model, which states that owners of capital (e.g., ETH) need to be compensated for taking on more perceived risk. Interestingly, neither the coefficient on the ETH price crash dummy and nor that of FTX collapse is significant (0.0186 and 0.0349, respectively).
Examining the results for solo stakers, the null hypothesis of no correlation could not be rejected for the dollar rewards rate (-0.0308), though, again the coefficient on risk free rate was positive and actually higher than for the overall population (0.1015 vs. 0.0714). Again the coefficient on the FTX collapse was not significant. More interestingly, the effect from ETH’s flash crash was highly negative and significant (-0.1434) at 0.1%, indicating that while the overall population did not discernibly react to the crash, solo stakers did, reducing the amount staked by over 14%.
A possible way of interpreting these results is that solos are motivated by a sense of altruism and are not generally swayed by small changes in dollar rewards, yet due to their limited ability to benefit from economies of scale and leverage, they generally have a higher minimal rewards threshold than non-solo stakers, below which solo stakers exit the market. Simply put, the prospect of markedly reduced future earnings causes solos to react more stronger than the overall population would, hurting decentralization. This is in line with the theoretical model’s finding that solos are more sensitive to changes in rewards than other types of stakers.
Table 8: 2SLS with event dummies
OLS 2SLS 2SLS Solos
ETH Rewards $0.2086^{***}$ $0.4640^{***}$ $-0.0308$
$(0.023)$ $(0.042)$ $(0.019)$
ETH Flash Crash $0.1753^{***}$ $0.0186$ $-0.1434^{***}$
$(0.021)$ $(0.031)$ $(0.014)$
FTX Collapse $-0.1085^{***}$ $0.0349$ $-0.0246$
$(0.033)$ $(0.042)$ $(0.019)$
Risk-free rate $0.2190^{***}$ 0.07140.0714{**}0.0714 ∗ ∗ $0.1015^{***}$
$(0.020)$ $(0.029)$ $(0.013)$
Constant $7.9750^{***}$ $3.1535^{***}$ $10.1725^{***}$
$(0.443)$ $(0.795)$ $(0.358)$
Observations 423 423 423
F-statistic 655.4 524.5 189.0
R-squared $0.862$ $0.823$ $0.649$
Standard errors in parentheses
*** p$<$0.01, ** p$<$0.05, * p$<$0.1
5.3.2 Gas fees as an IV
In our second estimation approach, we perform separate regressions for solo staking and for total staking, with results shown in Tables 9 and 10, respectively. There are two notable insights from these results. In the first-stage regressions, gas fees exhibit a positive and statistically significant effect on staking rewards, with F-statistics exceeding the threshold of 10, thereby meeting the relevance condition for instruments. This strong association reinforces the appropriateness of gas fees as an instrumental variable for staking rewards. In the second stage, the estimated elasticity of staking supply with respect to staking rewards is 1.184 for solo stakers alone and 1.078 for all stakers, both statistically significant at the 1 percent level. This result suggests that, consistent with Eloranta and Helminen, (2025), solo stakers are slightly more responsive to changes in staking rewards than the overall staking cohort.
We validate the robustness of these findings through additional tests. Following a similar approach to Cong et al., (2022), we incorporate various lags for both staking rewards and gas fees, observing minimal variation in results, which supports the stability of our estimates. Furthermore, when aggregating the data to a weekly frequency (see Tables 11 and 12 in the Appendix), the results remain consistent, underscoring the reliability of our elasticity estimates.
Again, the results are broadly consistent with the results of the theoretical model, which suggest that solo stakers are more sensitive to changes in the issuance yield compared to other staking categories. Still, these estimates are based on very short-term variations in gas fees. As such, it may be difficult to extrapolate the estimation results to an equilibrium model, such as the one developed above.
Table 9: Solo Staking, Daily Data
OLS 1st Stage 2SLS
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Log Total Staked Log Total Staked Log Total Staked Log Rewards (USD) Log Rewards (USD) Log Rewards (USD) Log Total Staked Log Total Staked Log Total Staked
(USD) (USD) (USD) 1-day Lag 7-day Lag (USD) (USD) (USD)
Log Rewards (USD) $0.678^{***}$ $1.184^{***}$
$(0.067)$ $(0.073)$
Log Rewards (USD), 1-day Lag $0.658^{***}$ $1.176^{***}$
$(0.071)$ $(0.074)$
Log Rewards (USD), 7-day Lag $0.587^{***}$ $1.146^{***}$
$(0.071)$ $(0.078)$
Log Gas Fees (USD) $0.228^{***}$
$(0.009)$
Log Gas Fees (USD), 1-day Lag $0.227^{***}$
$(0.009)$
Log Gas Fees (USD), 7-day Lag $0.228^{***}$
$(0.009)$
Constant $12.835^{***}$ $13.086^{***}$ $13.931^{***}$ $8.464^{***}$ $8.471^{***}$ $8.458^{***}$ $6.774^{***}$ $6.868^{***}$ $7.240^{***}$
$(0.797)$ $(0.853)$ $(0.853)$ $(0.143)$ $(0.143)$ $(0.142)$ $(0.877)$ $(0.888)$ $(0.937)$
Observations 622 621 615 622 621 615 622 621 615
R-squared $0.287$ $0.268$ $0.217$ $0.499$ $0.500$ $0.505$ $0.128$ $0.101$ $0.021$
Robust standard errors in parentheses
*** p$<$0.01, ** p$<$0.05, * p$<$0.1
Table 10: Total Staking, Daily Data
OLS 1st Stage 2SLS
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Log Total Staked Log Total Staked Log Total Staked Log Rewards (USD) Log Rewards (USD) Log Rewards (USD) Log Total Staked Log Total Staked Log Total Staked
(USD) (USD) (USD) 1-day Lag 7-day Lag (USD) (USD) (USD)
Log Rewards (USD) $1.279^{***}$ $1.078^{***}$
$(0.018)$ $(0.035)$
Log Rewards (USD), 1-day Lag $1.276^{***}$ $1.075^{***}$
$(0.019)$ $(0.036)$
Log Rewards (USD), 7-day Lag $1.269^{***}$ $1.069^{***}$
$(0.022)$ $(0.039)$
Log Gas Fees (USD) $0.406^{***}$
$(0.019)$
Log Gas Fees (USD), 1-day Lag $0.405^{***}$
$(0.019)$
Log Gas Fees (USD), 7-day Lag $0.407^{***}$
$(0.019)$
Constant $4.618^{***}$ $4.668^{***}$ $4.804^{***}$ $9.212^{***}$ $9.226^{***}$ $9.190^{***}$ $7.739^{***}$ $7.786^{***}$ $7.900^{***}$
$(0.283)$ $(0.298)$ $(0.338)$ $(0.292)$ $(0.292)$ $(0.290)$ $(0.543)$ $(0.556)$ $(0.605)$
Observations 622 621 615 622 621 615 622 621 615
R-squared $0.880$ $0.872$ $0.846$ $0.434$ $0.434$ $0.453$ $0.858$ $0.851$ $0.825$
Robust standard errors in parentheses
*** p$<$0.01, ** p$<$0.05, * p$<$0.1
Table 11: Solo Staking, Weekly Data
OLS 1st Stage 2SLS
(1) (2) (3) (4) (5) (6)
Log Total Staked Log Total Staked Log Rewards (USD) Log Rewards (USD) Log Total Staked Log Total Staked
(USD) (USD) 1-week Lag (USD) (USD)
Log Rewards (USD) $0.900^{***}$ $1.222^{***}$
$(0.171)$ $(0.166)$
Log Rewards (USD), 1-week Lag $0.791^{***}$ $1.184^{***}$
$(0.181)$ $(0.178)$
Log Gas Fees (USD) $0.246^{***}$
$(0.015)$
Log Gas Fees (USD), 1-week Lag $0.246^{***}$
$(0.015)$
Constant $10.367^{***}$ $11.889^{***}$ $10.129^{***}$ $10.127^{***}$ $5.871^{**}$ $6.412^{***}$
$(2.376)$ $(2.520)$ $(0.230)$ $(0.229)$ $(2.313)$ $(2.479)$
Observations 88 87 88 87 88 87
R-squared $0.355$ $0.281$ $0.751$ $0.755$ $0.309$ $0.212$
Robust standard errors in parentheses
*** p$<$0.01, ** p$<$0.05, * p$<$0.1
Table 12: Total Staking, Weekly Data
OLS 1st Stage 2SLS
(1) (2) (3) (4) (5) (6)
Log Total Staked Log Total Staked Log Rewards (USD) Log Rewards (USD) Log Total Staked Log Total Staked
(USD) (USD) 1-week Lag (USD) (USD)
Log Rewards (USD) $1.338^{***}$ $1.085^{***}$
$(0.028)$ $(0.081)$
Log Rewards (USD), 1-week Lag $1.327^{***}$ $1.078^{***}$
$(0.039)$ $(0.090)$
Log Gas Fees (USD) $0.452^{***}$
$(0.048)$
Log Gas Fees (USD), 1-week Lag $0.453^{***}$
$(0.048)$
Constant $3.041^{***}$ $3.257^{***}$ $10.439^{***}$ $10.428^{***}$ $7.468^{***}$ $7.610^{***}$
$(0.482)$ $(0.675)$ $(0.753)$ $(0.753)$ $(1.404)$ $(1.563)$
Observations 88 87 88 87 88 87
R-squared $0.913$ $0.883$ $0.506$ $0.528$ $0.880$ $0.852$
Robust standard errors in parentheses
*** p$<$0.01, ** p$<$0.05, * p$<$0.1
6 Discussion